OpenStudy (anonymous):

PLEASE HELP! What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x - 3 y = 5x - 1 y = 5x + 3 y = 5x - 19

OpenStudy (anonymous):

last

OpenStudy (anonymous):

How do you get that?

OpenStudy (anonymous):

y=mx+n but slope =5 so m=5 y=4x+n 1=4*5+n n=1-20=-19 y=5x-19

OpenStudy (anonymous):

thank you!